OpenStudy (anonymous):

Solve the following equation. One point is given for showing steps and one point is given for the correct solution. (x-3) / x = (x-3) / (x-6)

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

multiply thru by the lcm x(x-6):- (x-3)(x-6) = x(x-3) x^2 - 9x + 18 = x^2 - 3x -6x = -18 x = 3